# Positional Encoding Helps Recurrent Neural Networks Handle a Large Vocabulary

### v3

@@ 4.3 Metric of Gradient Stability
This study assessed the gradient stability of RNNs by feeding them pairs of input sequences that shared the initial target token ($t=1$) and varied in the subsequent disturbant tokens ($2\leq t\leq L$; see Fig. 4). Each pair of input sequences, denoted as $A$ and $B$, established two distinct, but ideally similar mappings, $f^{(A)}$ and $f^{(B)}$, from the first to the last latent state of the RNNs ($\vec{\mathbf{h}}_{2L}^{(s)}=f^{(s)}(\vec{\mathbf{z}}_{1})$, where $s\in\{A,B\}$). The gradient stability of the RNNs was defined by the dot-product similarities between the normalized gradients of these paired mappings:
@@
$\textsc{Stability}(A,B):=$ $\sum_{i=1}^{D}\langle\alpha_{i}^{(A)}\nabla f_{i}^{(A)}(\vec{\mathbf{z}}_{1}),\alpha_{i}^{(B)}\nabla f_{i}^{(B)}(\vec{\mathbf{z}}_{1})\rangle$
$=$ $\sum_{i=1}^{D}\alpha_{i}^{(A)}\alpha_{i}^{(B)}\sum_{j=1}^{(2)D}\frac{\partial h_{2L,i}^{(A)}}{\partial z_{1,j}}\cdot\frac{\partial h_{2L,i}^{(B)}}{\partial z_{1,j}}$ (3)
@@
$\alpha_{i}^{(s)}:=$ $\frac{\|\nabla f_{i}^{(s)}(\vec{\mathbf{z}}_{1})\|}{\sum_{k=1}^{D}\|\nabla f_{k}^{(s)}(\vec{\mathbf{z}}_{1})\|}$
$=$ $\left.\left(\sqrt{\sum_{j=1}^{(2)D}\Biggl{\lvert}\frac{\partial h_{2L,i}^{(s)}}{\partial z_{1,j}}\Biggr{\rvert}^{2}}\right)\middle/\left(\sum_{k=1}^{D}\sqrt{\sum_{j=1}^{(2)D}\Biggl{\lvert}\frac{\partial h_{2L,k}^{(s)}}{\partial z_{1,j}}\Biggr{\rvert}^{2}}\right)\right.$ (4)
@@
Acknowledgements
@@
This study was supported by JST ACT-X (JPMJAX21AN) and Core Research for Evolutional Science and Technology (JPMJCR17A4); JSPS Grant-in-Aid for Early-Career Scientists (JP21K17805) and for Scientific Research A (JP24H00774), B (JP22H03914), and C (JP24K15087); and Kayamori Foundation of Informational Science Advancement (K35XXVIII620). The author also gratefully acknowledges the support of the ACCMS, Kyoto University, regarding the use of their supercomputer system.
@@ A.1 Reverse-Ordering + Delayed-Addition
This task was too challenging to GRUs—even after reducing the input length to $L=16$—so only the results from LSTMs are reported below. Also, the network was trained for 600,000 iterations (i.e., twice longer than the other tasks) for ensuring the convergence. The other conditions/hyperparameters were the same as reported in the Materials and Methods section of the main text.
@@ Appendix B Robustness to Variations in Input Length
Appendix C Alternative implementations of Positional Encoding
@@
[…]
Figure 8: Token-wise accuracy of the sorting task performed by RNNs with and without positional encoding (labeled as “Position-Encoded” and “Vanilla” respectively). The input length was fixed at $L:=64$. The error bars represent the 95% confidence interval estimated from 10,000 bootstrapped samples of five training-test trials with different random seeds. Each of the five trials held out 1024 random sequences (= 65,536 tokens) for computing the test accuracy.

### v5

@@ 4.3 Analysis of Gradient Stability
Formally, the paired input sequences, denoted as $A$ and $B$, established two distinct, but ideally similar mappings, $f^{(A)}$ and $f^{(B)}$, from the first to the last latent state of the RNNs ($\vec{\mathbf{h}}_{2L}^{(s)}=f^{(s)}(\vec{\mathbf{z}}_{1})$, where $s\in\{A,B\}$). The gradient stability of the RNNs was defined by the dot-product similarities between the normalized gradients of these paired mappings:
@@
$\textsc{Stability}(A,B):=\sum_{i=1}^{D}\langle\alpha_{i}^{(A)}\nabla f_{i}^{(A)}(\vec{\mathbf{z}}_{1}),\alpha_{i}^{(B)}\nabla f_{i}^{(B)}(\vec{\mathbf{z}}_{1})\rangle=\sum_{i=1}^{D}\alpha_{i}^{(A)}\alpha_{i}^{(B)}\sum_{j=1}^{(2)D}\frac{\partial h_{2L,i}^{(A)}}{\partial z_{1,j}}\cdot\frac{\partial h_{2L,i}^{(B)}}{\partial z_{1,j}}$ (3)
@@
$\alpha_{i}^{(s)}:=\frac{\|\nabla f_{i}^{(s)}(\vec{\mathbf{z}}_{1})\|}{\sum_{k=1}^{D}\|\nabla f_{k}^{(s)}(\vec{\mathbf{z}}_{1})\|}=\left.\left(\sqrt{\sum_{j=1}^{(2)D}\Biggl{\lvert}\frac{\partial h_{2L,i}^{(s)}}{\partial z_{1,j}}\Biggr{\rvert}^{2}}\right)\middle/\left(\sum_{k=1}^{D}\sqrt{\sum_{j=1}^{(2)D}\Biggl{\lvert}\frac{\partial h_{2L,k}^{(s)}}{\partial z_{1,j}}\Biggr{\rvert}^{2}}\right)\right.$ (4)
@@
Unlike the RNN models, both the vanilla and position-encoded S4Ds achieved high accuracy over 96% for the initial target token ($t=1$), regardless of the frequency of the target and disturbants. Accordingly, for the analysis of S4D, the target token was positioned in the middle at $t=23$, where the vanilla model exhibited its poorest accuracy with the Rare disturbants. The disturbants were prefixed and suffixed to this target to construct input sequences. The prefix disturbants were shared between the paired sequences, thereby ensuring that the latent dynamics of the model remained identical up to the target token.
It should also be noted that the latent states of S4D are complex-valued (while its outputs are real-valued), and consequently, the gradients and their dot-product similarities are also complex-valued. For the sake of the present analysis, the complex-valued gradients were treated as the double-sized real arrays, and a real-valued similarity was defined by Eq. 3. This is equivalent to taking the real component of the complex-valued similarity, and is intuitively natural given that a perfect alignment between complex gradient directions yields a real-valued score of 1.0 (= the norm of a normalized complex vector). Additionally, the extra dimension in the latent states representing the order of the Legendre polynomials was merged with the channel dimension, and the entire state was treated as a flattened vector.
Figure 5: Gradient stability of GRU/LSTM/S4D trained on the reverse-ordering task with and without positional encoding (labeled as “Position-Encoded” and “Vanilla” respectively). For the GRU and LSTM, the stability was defined by the dot-product similarity of latent-to-latent gradients after normalization over the output dimensions, conditioned on two input sequences sharing the initial “target” token (whose Frequent vs. Rare distinction is represented by the line color), followed by Frequent or Rare disturbants (represented by the solid vs. dashed lines). For the S4D, the target token was positioned at $t=23$, where the vanilla model scored the worst accuracy with the Rare disturbants. The disturbants were prefixed and suffixed to the target to construct input sequences. The prefix disturbants were shared between the paired sequences, ensuring that the latent dynamics of the model was guaranteed to remain identical up to the target token. The total input length was $1+63=22+1+41=64$. The average similarity over 1024 input pairs times five trials is plotted for every 5000 training iterations.
By contrast, the impact of positional encoding on the gradient stability of the S4D was marginal; unlike the RNNs, the vanilla S4D was highly stable by itself against Rare disturbants throughout the training, even though there was a visible relative destabilization due to Rare disturbants compared to Frequent disturbants in the early stages of training, as well as an observable improvement by positional encoding. It is also noteworthy that the difference between the Frequent vs. Rare disturbants diminished after 10,000 training iterations. Consequently, gradient stability does not fully account for the decline in the accuracy of S4D in the presence of Rare disturbants, nor does it explain the enhancement brought about by positional encoding.
5 Discussion
5.1 Difficulties in Handling a Large Vocabulary
This study examined the RNN gradients and identified their destabilization when processing low-frequency tokens, which are necessarily included in a large vocabulary. Specifically, inputs that do not contribute to gradient-based optimization at a target time step (e.g., tokens at $2\leq t\leq L$ upon the retrieval of the initial token at $t=2L$ in the reverse-ordering task) were found to be detrimental.
In general cases of time series processing, data points carrying crucial information for specific time steps become irrelevant otherwise. Consequently, each token exhibits a dual nature—both crucial and noisy—throughout the task, and processing rare tokens is particularly challenging presumably because they are irrelevant at most of the time while making a large impact on the learning through the greater loss to compensate for their fewer learning opportunities. Dealing with such “unignorable noise” presents a pervasive challenge for RNNs.
Although low-frequency tokens destabilize the gradient-based learning of RNNs, the present study also discovered that this issue can be alleviated by positional encoding. This enhancement of RNNs via positional encoding is noteworthy because RNNs were specifically designed to process time series data on their own; hence, unlike Transformers, they are presumed to function without relying on an “external clock” (Siegelmann and Sontag,, 1992, 1994, 1995; Siegelmann,, 1996, 1999). Consequently, position-encoded RNNs have remained largely unexplored, with only two exceptions to the best of the author’s knowledge (Vincent-Lamarre et al.,, 2016; Karanikolos and Refanidis,, 2019). The findings of the present study—namely, the improvement in the manageable vocabulary size due to the enhanced gradient stability—broaden the currently limited understanding of the impact of positional encoding on RNNs.
5.3 Limitations and Future Directions
A primary unresolved question in this study pertains to the mechanism behind the gradient stabilization by positional encoding. All the findings here are based on experimental investigations, lacking rigorous mathematical explanations for how and why the gradients of RNNs are destabilized by infrequent tokens and stabilized by positional encoding. Moreover, the present study primarily focused on the canonical implementation of sinusoidal positional encoding designed for Transformers (Eqs. 1,2), leaving it open which parameters of the sinusoidal waves (i.e., frequencies and phases) are critical for gradient stabilization. Future research may broaden its scope to encompass more general forms of positional encoding, such as wavelets and non-periodic signals (see Appendix D for a preliminary comparison among sinusoidal, learnable, and randomly-fixed encodings).
Moreover, the analysis of gradient stability did not fully address the enhanced performance of the position-encoded state-space model (S4D). In terms of accuracy, the positioned-encoded S4D exhibited greater robustness to infrequent tokens compared to the vanilla model, resembling the behavior observed in RNNs. However, the gradients of the vanilla S4D were too stable to account for this descline in performance. This leaves the question open how positional encoding influences gradient-based learning of state-space models. Additionally, future studies may investigate a broader range of state-space models—including the state-of-the-art architecture of Mamba (Gu and Dao,, 2024)—to achieve a comprehensive understanding of the interplay between positional encoding and these models.
In addition to these scientifically oriented questions, future studies could also address practical applications of position-encoded RNNs and neural state-space models. Although positional encoding enhanced model performance across different synthetic tasks (see Appendix A), the extent of this enhancement is task-dependent. Indeed, while Karanikolos and Refanidis, (2019) reported the effectiveness of positional encoding for an LSTM text summarizer, the present study found no empirical advantage for the language modeling task, aside from a slightly more rapid decline in training loss (see Appendix E). Thus, positional encoding is not a panacea for arbitrary tasks, and further investigations are necessarily to determine when it is effective.
Acknowledgments
@@
This study was supported by JST ACT-X (JPMJAX21AN) and Core Research for Evolutional Science and Technology (JPMJCR17A4, JPMJCR22P5); JSPS Grant-in-Aid for Early-Career Scientists (JP21K17805) and for Scientific Research A (JP24H00774), B (JP22H03914), and C (JP24K15087); and Kayamori Foundation of Informational Science Advancement (K35XXVIII620). The author also gratefully acknowledges the support of the ACCMS, Kyoto University, regarding the use of their supercomputer system.
@@ A.1 Reverse-Ordering + Delayed-Addition
This task was too challenging to GRUs—even after reducing the input length to $L=16$—so only the results from LSTMs are reported below. Also, the network was trained for 600,000 iterations (i.e., twice longer than the other tasks) for ensuring the convergence. The other conditions/hyperparameters were the same as reported in the main text.
@@ A.2 Sorting
Figure 8: Token-wise accuracy (left) and sequence-wise reconstruction errors (right) of the sorting task performed by RNNs with and without positional encoding (labeled as “Position-Encoded” and “Vanilla” respectively). The input length was fixed at $L:=64$. The error bars represent the 95% confidence interval estimated from 10,000 bootstrapped samples of five training-test trials with different random seeds. Each of the five trials held out 1024 random sequences (= 65,536 tokens) for computing the test accuracy.
@@
A.3 Predecessor Query
[…]
Figure 9: Illustration of the precedent-query task.
Finally, this section presents benchmark results for the predecessor-query task, illustrated in Fig. 9. The network first received a sequence of non-repeating random integers, $x_{1},\dots,x_{L}$. Subsequently, one of the non-initial input integers, $x_{t_{\textrm{query}}}(2\leq t_{\textrm{query}}\leq L)$, was randomly selected and reintroduced to the network at time $t=L+1$. The learning objective is to return the predecessor of the reviewed integer ($=x_{t_{\textrm{query}}-1}$). The predecessor-query task evaluates the capacity of RNNs to integrate information regarding both the order and content of input sequences.
As in the reverse-ordering + delayed-addition task, the input sequence was reduced to $L=16$ due to the complexity of the task, and the experiment focused on the LSTM. The number of training iterations was maintained at 300,000.
Figure 10: Accuracy of the precedent-query task performed by the LSTM with and without positional encoding (labeled as “Position-Encoded” and “Vanilla” respectively). The input length was fixed at $L:=16$. The error bars represent the 95% confidence interval estimated from 10,000 bootstrapped samples of five training-test trials with different random seeds. Each of the five trials held out 1024 random sequences (= 65,536 tokens) for computing the test accuracy.
Similar to the other benchmarks, positional encoding improved the LSTM’s capacity to manage the larger vocabularies.
@@ Appendix B Robustness to Variations in Input Length
Appendix C Effects of Additional Parameters in Position-Encoded RNNs
The concatenation of positional encoding with input embeddings inflates the number of learnable parameters in the input-to-hidden projection weights. This additional parameterization per se does not influence the learning of the input embeddings, and therefore does not elucidate the enhanced performance of position-encoded RNNs. This section substantiates this argument by equalizing the number of learnable parameters between the vanilla and position-encoded models.
Specifically, the equalization was achieved by concatenating two identical copies of the input embeddings and feeding them to the LSTM. This configuration—henceforth termed “double vanilla”—effectively doubled the size of the input-to-hidden weight for each gate in the LSTM, aligning it with that of the position-encoded LSTM, while maintaining all other parameters, including the dimensionality of the (non-repeated) input embeddings.
Figure 12: Token-wise accuracy of the reverse-ordering and sorting tasks performed by the LSTM with and without positional encoding. The “double vanilla” model concatenated two copies of the identical input embeddings to match the number of parameters with the position-encoded model. The input length was fixed at $L:=64$. The vocabulary size was set to 16,384. The error bars represent the 95% confidence interval estimated from 10,000 bootstrapped samples of five training-test trials with different random seeds. Each of the five trials held out 1024 random sequences (= 65,536 tokens) for computing the test accuracy.
As illustrated in Fig. 12, the double vanilla LSTM did not yield any improvements in the reverse-ordering or sorting tasks. These results affirm that the reported enhancement of RNNs is not merely attributable to the additional parameterization associated with the positional encoding.
Appendix D Alternative Implementations of Positional Encoding
@@
Appendix E Language Modeling
This section reports benchmark results for the language modeling task. Single-layer LSTMs with and without sinusoidal positional encoding were trained and tested on the WikiText-103 dataset (Merity et al.,, 2017). Due to constraints in computational resources, the vocabulary was reduced from the original size of 267,735 to 32,768 by retokenizing the raw data using SentencePiece (Kudo and Richardson,, 2018). The headings were removed, and the main text was segmented by paragraphs (separated by the line break). Additionally, only the first 1024 tokens of each paragraph were utilized for training and testing, ensuring that the absolute positional encoding always aligned with the beginning of each paragraph. The hyperparameters were configured as specified in §3.3.
Figure 15: Training perplexities on the WikiText-103 dataset. Each of the five trials with a different random seed is represented by a single line.
Table 1: Test perplexities on the WikiText-103 dataset. The minimum, mean, and maximum are obtained from five trials with different random seeds.
Model Min Mean Max
Vanilla LSTM 36.8257 37.7731 38.916589
Position-Encoded LSTM 38.0685 38.5384 38.893656
As illustrated in Fig. 15, positional encoding proved effective only for marginally faster learning during the initial phase of training. The difference diminished around 10,000/30,000 iterations, and the test perplexities of the position-encoded model were inferior to those of the vanilla model.